OpenStudy (anonymous):

A floor has two square-shaped designs. The area of the second square-shaped design is sixteen times greater than the area of the first square-shaped design. Which statement gives the correct relationship between the lengths of the sides of the two squares?

OpenStudy (anonymous):

The length of the side of the second square is 16 times greater than the length of the side of the first square. The length of the side of the second square is 2 times greater than the length of the side of the first square. The length of the side of the second square is 4 times greater than the length of the side of the first square. The length of the side of the second square is 8 times greater than the length of the side of the first square.

OpenStudy (campbell_st):

look at it this way let the area of the 2nd square be 16 the formula for area of a square is \[A = l^2 \] l = side length so for the 2nd square 16 = l^2 to find the side length take the square root of both sides let the 1st side have area 1 then \[1 = l^2 \] solve for l by taking the square root of both sides... now you can compare the side lengths of square 2 and square 1